OpenStudy (anonymous):

How far from the origin is the point (-5, -12)? A. 13 B. 7 C. 11.49 D. 84.5

OpenStudy (anonymous):

what formula do i use

OpenStudy (anonymous):

d=\[\sqrt{(-5-0)^2+(-12-0)^2}=\sqrt{169}=13\]

OpenStudy (anonymous):

\[D = \sqrt{(x2-x1)^2 + (y2-y1)^2}\] the two points are the origin (0,0) and the other one (-5,-12) substitute them in the equation.

OpenStudy (anonymous):

\[d=\sqrt{(x _{2}-x _{1})^2+(y _{2}-y _{1})^2}\]

OpenStudy (anonymous):

soo if theres one equation in this situation i make the other equation (0,0)

OpenStudy (anonymous):

13 is the ans. its ans can be find out by distance formula as ASADD123 stated.

OpenStudy (anonymous):

ohhh o